# Neuromorphic Computing with Microfluidic Memristors

### v1

@@ Abstract
Conical microfluidic channels filled with electrolytes exhibit volatile memristive behavior, offering a promising foundation for energy-efficient, neuromorphic computing. Here, we integrate these iontronic channels as additional nonlinear element in nonlinear Shinriki-inspired oscillators and demonstrate that they exhibit alternating chaotic and non-chaotic dynamics across a broad frequency range. Exploiting this behavior, we construct XOR and NAND gates by coupling three “Memriki” oscillators, and we further realize the full set of standard logic gates through combinations of NAND gates. Our results establish a new paradigm for iontronic computing and open avenues for scalable, low-power logical operations in microfluidic and bio-inspired systems.
@@ I Introduction
The era we live in has been proposed the Silicon age [1], defined by semiconductor transistors that have revolutionized data processing, accelerating it by several orders of magnitude. This technological advancement has enabled groundbreaking inventions, including computers, the internet, and, more recently, artificial intelligence (AI). However, advancements in AI are rapidly approaching a critical power bottleneck [2]. A key factor driving this challenge is the high energy cost caused by Ohmic losses of frequent data transfer between physically separated processing and storage units in conventional computing architectures, a limitation known as the von Neumann bottleneck [3]. AI power consumption is currently doubling every four to six months [4], which would surpass global energy production by the 2030s. This trend highlights the pressing need for novel computing paradigms. To enhance energy efficiency and meet the increasing demands for computing power in the era of AI and big data, a fundamentally different approach is required.
@@
Inspired by the most energy-efficient computer known – the biological brain – the field of neuromorphic computing emerged [5]. The human brain, an intricate biological neural network comprising approximately 1011 neurons and 1015 synapses, operates with a remarkable efficiency at just $20\,\mathrm{W}$ of power. A key subfield of neuromorphic computing focuses on developing devices that mimic the computational functions of the brain. One such device is the memristor, first theoretically proposed by Chua in 1971 as the fourth basic circuit element [6]. Memristors act as memory elements, with a resistance that depends on past voltage or current, allowing them to retain information. In 2008, Strukov et al. experimentally demonstrated a solid-state memristor, paving the way for its use as an artificial synapse in neuromorphic systems [7], with present-day applications in for instance crossbar arrays [8, 9]. However, silicon-based technologies typically rely on a single type of information carrier: electrons (or their absence (holes)). In contrast, the biological brain utilizes ions dissolved in water as information carriers, allowing for a diverse range of soluble information carriers, such as various species of ions and small molecules. A key distinction between solid-state electronics and ion-mediated biological systems lies in their energy efficiency, with biological systems operating orders of magnitude more efficiently, albeit at much lower clock frequencies. This contrast has led to the emergence of iontronics, a subfield of neuromorphic computing focused on developing electronic circuit elements where ions dissolved in water serve as information carriers [10]. Just as solid-state memristors have been widely explored, researchers have pursued iontronic analogs [11]. Memristive effects in charged conical nanopores were first observed in 2010 and 2012 [12, 13], paving the way for the development of diverse nanofluidic memristors based on various mechanisms. These include electric double layer polarization [13, 14, 15, 16], salinity gradients yielding negative differential resistance behavior [17], electro-wetting [18, 19], structural and conformational changes of nanopores [20, 21], the ionic analogue [22] of electronic Coulomb blockade [23], mixtures of and interfaces between ionic liquids and water [24, 25], and specific polyelectrolyte-ion interactions [26].
@@
The next step towards neuromorphic computing is integrating iontronic devices into electric circuits. Key questions include: Can these networks emulate neural functions, perform Boolean logic operations, or execute simple machine learning tasks [27]? While the use of iontronic elements in large-scale networks is still in its infancy, initial progress has been made in implementing basic logic gates such as AND, OR [28], NAND and NOR [29], and material implication IMP [15] gates. However, a complete set of logic gates based on (memristive) iontronics has yet to be realized.
@@
We first provide a brief introduction to volatile microfluidic memristors as descibeed in Ref. [30], and then explore their integration into electric circuits that form Shinriki-inspired (non-linear) oscillators [31]. Finally, we demonstrate how these oscillators can be combined to realize the full set of logic gates, thereby advancing the use of iontronics for neuromorphic and Boolean logic applications.
@@ II Microfluidic iontronic memristor
Figure 1: (a) Schematic of a conical channel filled with an electrolyte (cations in green, anions in red) with a length $L$, base radius $R_{b}$, tip radius $R_{t}$, bulk ion densities $\rho_{\pm}$, and surface charge $e\sigma$. A time-dependent voltage $U(t)$ is applied across the channel. (b) Limit cycles of the conductance-voltage relation for three different driving frequencies $f=0.4\,\mathrm{Hz}$ (blue), $f=40\,\mathrm{Hz}$ (orange), and $f=4000\,\mathrm{Hz}$ (green) of a sinusoidal potential with an amplitude of 1 Volt. (c) Limit cycles of the current-voltage relation for the same frequencies and parameters as in (b).
In this study, we focus on volatile microfluidic memristors as presented in Ref. [30] and illustrated in Fig. 1(a). This conical channel, with a length $L$, has a wide opening at $x=0$ and a narrow tip at $x=L$, and connects two reservoirs of an aqueous 1:1 electrolyte at a total bulk ion concentration $2\rho_{b}$ at room temperature. The channel can carry an ionic electric current when driven by an applied potential difference between the two reservoirs. The combination of the conical channel geometry and a negative surface charge $\sigma$ on the channel walls results in current rectification when a static voltage drop $U$ is applied [34]. This phenomenon was theoretically explained in terms of a nontrivial $U$-dependent steady-state salt concentration profile $\rho_{s}(x;U)$ for $x\in[0,L]$ that build up within the channel [35]. This concentration profile is caused by the much stronger electric field near the tip compared to the base and leads to ionic depletion or accumulation depending on the polarity of the potential. Consequently, the steady-state electric conductance $g_{\infty}(U)$ depends non-trivially on $U$ and exhibits a hysteresis loop in the current-voltage relation when a time-dependent oscillating potential $U(t)$ is applied with a period on the order of the build-up time $\tau$ of the ionic concentration profile [30]. The time-dependent conductance $g(t)$ of such a memristive channel was calculated from numerical solutions of the Poisson-Nernst-Planck-Stokes (partial differential) equations for the coupled transport of charge, water, and ions [30]. In a large parameter regime $g(t)$ was also found to be accurately approximated by the following (ordinary differential) equation of motion 111this is only an approximation of the steady-state conductance, but Kamsma et al.  found that it is reasonable within the applied voltage range [30, 44]
@@ III Characterization of a Single Memriki oscillator
To leverage the volatile behavior of the memristors of Section II, we incorporate two of them into a nonlinear electrical circuit with an intrinsic oscillatory timescale similar to that of the memristors. The circuit design, shown in Fig. 2, is heavily inspired by the well-known Shinriki-oscillator [31], albeit with a slightly modified topology and with iontronic memristors $M_{1}$ and $M_{2}$ rather than diodes. We propose the name Memriki oscillator for this circuit. The circuit, driven by a sinusoidal source potential $U_{in}$ with frequency $f$, consists of a negative differential resistance $\mathrm{NDR}$, two anti-parallel memristors in series with a (weakly damped) LRC oscillator with eigen frequency $1/(2\pi\sqrt{L_{o}C_{o}})$, and a resistor $R_{1}$ and capacitor $C_{1}$ that are both parallel to the memristor-oscillator part. Based on the original work of Shinriki et al.  [31], the $\mathrm{NDR}$ is an active element with a characteristic function given by $R(U)\,\mathrm{g_{0}}=(U/\,\mathrm{V})^{3}-3U/\,\mathrm{V}$. This implies that it behaves as an active amplifier rather than a passive resistor within the ±$\pm 1\,\mathrm{V}$ voltage range around $0\,\mathrm{V}$. The parallel resistor and capacitor, with $C_{1}=10^{-4}\,\mathrm{g_{0}s}$ and $g_{0}R_{1}$ of order unity, can serve as tunable parameters that significantly alter the dynamics of the circuit. The electric specifications for the oscillatory part of the circuit read $C_{o}=10^{-3}\,\mathrm{g_{0}s}$, $L_{o}=10^{-3}\,\mathrm{s}/\mathrm{g_{0}}$ and $R_{o}=10^{-3}\,\mathrm{g_{0}}^{-1}$, ensuring that $R_{o}\ll R_{1}$. As a result, the eigen period of the weakly damped oscillator is approximately $2\pi\sqrt{L_{o}C_{o}}\approx 6\,\mathrm{ms}$, which is comparable to the memory time of the two memristors, $\tau\approx 5\,\text{ms}$. While we describe and denote all circuit elements (except for the memristors) as conventional electrical components, we note that microfluidic counterparts exist for each of them [39, 40]. This enables a purely microfluidic realization of this Memriki oscillator.
The dynamic response of the Memriki oscillator varies dramatically depending on the precise values of the resistor $R_{1}$ and the driving frequency $f$. For instance, it can exhibit behaviors ranging from simple driven oscillations to chaos to subharmonic oscillations, as illustrated in  Fig. 3. A comprehensive study of the circuit’s full dynamics across a broader parameter space is beyond the scope of this paper. Instead, we focus here on the coupling of multiple Memriki oscillators.
Figure 2: Schematic of a single Memriki oscillator. The capacitor $C_{o}$, inductor $L_{o}$ and resistor $R_{o}$ together form a damped oscillator. The circuit is driven by the input voltage $U_{\mathrm{in}}$, measured relative to the reference voltage $U_{\mathrm{ref}}$. A negative differential resistance $\mathrm{NDR}$ introduces nonlinearity into the circuit. The output voltage $U_{\mathrm{out}}$ is measured, stabilized by $C_{1}$, and slowly relaxed towards $U_{\mathrm{ref}}$ by $R_{1}$. Two iontronic memristors $M_{1}$ and $M_{2}$ are added between $U_{\mathrm{out}}$ and the oscillatory section to introduce further nonlinearity.
@@
Figure 3: Dynamics of the driven Memriki oscillator from Fig. 2 for three different combinations of resistor $R_{1}$ and driving frequency $f$, with all other circuit elements specified in the text. The top panels (a)-(c) show the time series of the input voltage $U_{\mathrm{in}}=0.25\sin(2\pi ft)\,\mathrm{V}$ (orange) and the output voltage $U_{\mathrm{out}}$ (blue) after 100 periods. The bottom panels (d)-(f) show parametric plots of $U_{\mathrm{out}}$ versus $U_{\mathrm{in}}$ over 1000 periods, revealing stable periodic orbits in (d) and (f) separated by a chaotic, non-periodic regime in (e). Thus, in (a,d) the Memriki oscillator follows the input signal and oscillates at the driving frequency $f$. In (b,e), the response exhibits chaotic behavior, while in (c,f), the response is subharmonic, oscillating at a frequency of $f/3$.

### v2

@@ Abstract
Conical microfluidic channels filled with electrolytes exhibit volatile memristive behavior, offering a promising platform for energy-efficient, neuromorphic computing. Here, we integrate these iontronic channels as additional nonlinear elements in nonlinear Shinriki-inspired oscillators and demonstrate that they exhibit alternating chaotic and non-chaotic dynamics across a broad frequency range. Exploiting this behavior, we construct XOR and NAND gates by coupling three “Memriki” oscillators, and we further realize the full set of standard logic gates through combinations of NAND gates. Our results establish a new paradigm for iontronic computing and open avenues for scalable, low-power logical operations in microfluidic and bio-inspired systems.
@@ I Introduction
The era we live in has been proposed the Silicon age [1], defined by semiconductor transistors that have revolutionized data processing, accelerating it by several orders of magnitude. This technological advancement has enabled groundbreaking inventions, including computers, the internet, and, more recently, artificial intelligence (AI).
@@
However, advancements in AI are rapidly approaching a critical power bottleneck [2]. A key factor driving this challenge is the high energy cost caused by Ohmic losses of frequent data transfer between physically separated processing and information storage units in conventional computing architectures, a limitation known as the von Neumann bottleneck [3]. AI power consumption is currently doubling every four to six months [4], which would surpass global energy production by the 2030s. This trend highlights the pressing need for novel computing paradigms. To enhance energy efficiency and meet the increasing demands for computing power in the era of AI and big data, a fundamentally different approach is required [5].
@@
Inspired by the most energy-efficient computer known – the biological brain – the field of neuromorphic computing emerged [6]. The human brain, an intricate biological neural network comprising approximately 1011 neurons and 1015 synapses, operates with a remarkable efficiency at just $20\,\mathrm{W}$ of power. A key subfield of neuromorphic computing focuses on developing devices that mimic the computational functions of the brain. One such device is the memristor, first theoretically proposed by Chua in 1971 as the fourth basic circuit element [7]. Memristors act as memory elements, with a resistance that depends on past voltage or current, allowing them to retain information. In 2008, Strukov et al. experimentally demonstrated a solid-state memristor, paving the way for its use as an artificial synapse in neuromorphic systems [8], with present-day applications in for instance crossbar arrays [9, 10]. However, silicon-based technologies typically rely on a single type of information carrier: electrons (or their absence (holes)). In contrast, the biological brain utilizes ions dissolved in water as information carriers, allowing for a diverse range of soluble information carriers, such as various species of ions and small molecules. A key distinction between solid-state electronics and ion-mediated biological systems lies in their energy efficiency, with biological systems operating orders of magnitude more efficiently, albeit at much lower clock frequencies.
This contrast has led to the emergence of iontronics, a subfield of neuromorphic computing focused on developing electronic circuit elements where ions dissolved in water serve as information carriers [11]. Just as solid-state memristors have been widely explored, researchers have pursued iontronic analogs [12]. Memristive effects in charged conical nanopores were first observed in 2010 and 2012 [13, 14], paving the way for the development of diverse nanofluidic memristors based on various mechanisms. These include electric double layer polarization [14, 15, 16, 17], salinity gradients yielding negative differential resistance behavior [18], electro-wetting [19, 20], structural and conformational changes of nanopores [21, 22], the ionic analogue [23] of electronic Coulomb blockade [24], mixtures of and interfaces between ionic liquids and water [25, 26], and specific polyelectrolyte-ion interactions [27].
@@
The next step towards neuromorphic computing is integrating iontronic devices into electric circuits. Key questions include: Can these networks emulate neural functions, perform Boolean logic operations, or execute simple machine learning tasks [28]? While the use of iontronic elements in large-scale networks is still in its infancy, initial progress has been made in implementing basic logic gates such as AND, OR [29, 30], NAND and NOR [31], and material implication IMP [16] gates. A combination of currents and pH as inputs have been used to realize a set of logic gates [32, 33], requiring expensive pH changes, when switching inputs to the gate. Nanofluidic carbon nanotubes have been used to realize nanofluidic transistors which can be used for logic gates [34]. However, a complete set of logic gates based on (memristive) iontronics purely based on electronic inputs has yet to be realized.
@@
We first provide a brief introduction to volatile microfluidic memristors as described in Ref. [35], and then explore their integration into electric circuits that form Shinriki-inspired (non-linear) oscillators [36]. Finally, we demonstrate how these oscillators can be combined to realize the full set of logic gates, thereby advancing the use of iontronics for neuromorphic and Boolean logic applications.
@@ II Microfluidic iontronic memristor
Figure 1: (a) Schematic of a conical channel filled with an aqueous 1:1 electrolyte (cations in green, anions in red) with length $L$, base radius $R_{b}$, tip radius $R_{t}$, bulk ion concentrations $\rho_{\pm}=\rho_{b}$ in the reservoirs, and surface charge $e\sigma$. A time-dependent voltage $U(t)$ is applied across the channel. (b) Limit cycles of the conductance-voltage relation for three different driving frequencies $f=0.4\,\mathrm{Hz}$ (blue), $f=40\,\mathrm{Hz}$ (orange), and $f=4000\,\mathrm{Hz}$ (green) of a sinusoidal potential with an amplitude of 1 Volt. (c) Limit cycles of the current-voltage relation for the same frequencies and parameters as in (b).
In this study, we focus on volatile microfluidic memristors as presented in Ref. [35] and illustrated in Fig. 1(a). This conical channel of length $L$ has a wide opening of base radius $R_{b}$ at $x=0$ and a narrow tip of radius $R_{t}<R_{b}$ at $x=L$, and connects two reservoirs of an aqueous 1:1 electrolyte at a total bulk ion concentration $2\rho_{b}$ at room temperature. The channel can carry an ionic electric current when driven by an applied potential difference between the two reservoirs. The combination of the conical channel geometry and a negative surface charge $e\sigma$ on the channel walls results in current rectification when a static voltage drop $U$ is applied [39]. This phenomenon was theoretically explained in terms of a nontrivial $U$-dependent steady-state salt concentration profile $\rho_{s}(x;U)$ for $x\in[0,L]$ that build up within the channel [40]. This concentration profile is caused by the much stronger electric field near the tip compared to the base and leads to ionic depletion or accumulation depending on the polarity of the potential. Consequently, the steady-state electric conductance $g_{\infty}(U)$ depends non-trivially on $U$ and exhibits a hysteresis loop in the current-voltage relation when a time-dependent oscillating potential $U(t)$ is applied with a period on the order of the build-up time $\tau$ of the ionic concentration profile [35]. The time-dependent conductance $g(t)$ of such a memristive channel was calculated from numerical solutions of the Poisson-Nernst-Planck-Stokes (partial differential) equations for the coupled transport of charge, water, and ions [35]. In a large parameter regime $g(t)$ was also found to be accurately approximated by the following (ordinary differential) equation of motion 111this is only an approximation of the steady-state conductance, but Kamsma et al.  found that it is reasonable within the applied voltage range [35, 49]
@@
Figure 2: Schematic of a single Memriki oscillator. The capacitor $C_{o}$, inductor $L_{o}$ and resistor $R_{o}$ together form a damped oscillator. The circuit is driven by the input voltage $U_{\mathrm{in}}$, measured relative to the reference voltage $U_{\mathrm{ref}}$. A negative differential resistance $\mathrm{NDR}$ introduces nonlinearity into the circuit. The output voltage $U_{\mathrm{out}}$ is measured, stabilized by $C_{1}$, and slowly relaxed towards $U_{\mathrm{ref}}$ by $R_{1}$. Two iontronic memristors $M_{1}$ and $M_{2}$ are added between $U_{\mathrm{out}}$ and the oscillatory section to introduce further nonlinearity. With the switch SW1 the memristors can be replaced with ordinary resistors.
@@ III Characterization of a Single Memriki oscillator
Figure 3: Dynamics of the driven Memriki oscillator from Fig. 2 for three different combinations of resistor $R_{1}$ and driving frequency $f$, with all other circuit elements specified in the text. The top panels (a)-(c) show the time series of the input voltage $U_{\mathrm{in}}=0.25\sin(2\pi ft)\,\mathrm{V}$ (orange) and the output voltage $U_{\mathrm{out}}$ (blue) during 25 driving periods after 100 initial periods. The bottom panels (d)-(f) show the corresponding parametric plots of $U_{\mathrm{out}}$ versus $U_{\mathrm{in}}$ over 1000 driving periods, revealing stable periodic orbits in (d) and (f) separated by a chaotic, non-periodic regime in (e). Thus, in (a,d) the Memriki oscillator follows the input signal and oscillates at the driving frequency $f$, in (b,e) the response exhibits chaotic behavior, while in (c,f), the response is subharmonic with a threefold period at an oscillating frequency $f/3$.
To leverage the volatile behavior of the memristors of Section II, we incorporate two of them into a nonlinear electrical circuit with an intrinsic oscillatory timescale similar to that of the memristors. The circuit design, shown in Fig. 2, is heavily inspired by the well-known Shinriki-oscillator [36], albeit with a slightly modified topology and with iontronic memristors $M_{1}$ and $M_{2}$ rather than diodes. We propose the name Memriki oscillator for this circuit. Later we will swap out the memristors for conventional resistors to study the role of the memristors. In order to allow for this swapping, we introduced the switch SW1 into the circuit, to either connect the resistors or the memristors. The circuit, driven by a sinusoidal source potential $U_{in}=U_{0}\sin(2\pi ft)$ with amplitude $U_{0}$ and frequency $f$, consists of a negative differential resistance $\mathrm{NDR}$, two anti-parallel memristors in series with a (weakly damped) LRC oscillator with eigen frequency $1/(2\pi\sqrt{L_{o}C_{o}})$, and a resistor $R_{1}$ and capacitor $C_{1}$ that are both parallel to the memristor-oscillator part. Based on the original work of Shinriki et al.  [36], the $\mathrm{NDR}$ is an active element with a characteristic function given by $R(U)\,\mathrm{g_{0}}=(U/\,\mathrm{V})^{3}-3U/\,\mathrm{V}$. This implies that it behaves as an active amplifier rather than a passive resistor within the ±$\pm 1\,\mathrm{V}$ voltage range around $0\,\mathrm{V}$. The parallel resistor and capacitor, with $g_{0}R_{1}$ of order unity and $C_{1}=10^{-4}\,\mathrm{g_{0}s}$, respectively, can serve as tunable parameters that significantly alter the dynamics of the circuit. The electric specifications for the oscillatory part of the circuit read $C_{o}=10^{-3}\,\mathrm{g_{0}s}$, $L_{o}=10^{-3}\,\mathrm{s}/\mathrm{g_{0}}$ and $R_{o}=10^{-3}\,\mathrm{g_{0}}^{-1}$, ensuring that $R_{o}\ll R_{1}$. As a result, the eigen period of the weakly damped oscillator is approximately $2\pi\sqrt{L_{o}C_{o}}\approx 6\,\mathrm{ms}$, which is comparable to the memory time of the two memristors, $\tau\approx 5\,\text{ms}$. While we describe and denote all circuit elements (except for the memristors) as conventional electrical components, we note that microfluidic counterparts exist for each of them [44, 45]. This enables a purely microfluidic realization of this Memriki oscillator.
The dynamic response of the Memriki oscillator varies dramatically depending on the precise values of the resistor $R_{1}$ and the driving frequency $f$. For three different combinations of $R_{1}$ and $f$ (and for $U_{0}=0.25\,\mathrm{V}$ and all other parameters fixed as in the text above) this parameter sensitivity of the response is illustrated in Fig. 3. Here the top row (a)-(c) shows the time-dependence of the driving voltage $U_{\mathrm{in}}(t)=U_{0}\sin(2\pi ft)$ (orange) with the resulting output voltage $U_{\mathrm{out}}(t)$ (blue); the bottom row (d)-(e) shows the parametric plot $(U_{\mathrm{in}}(t),U_{\mathrm{out}}(t))$. Fig. 3 thus exhibits behaviors ranging from simple driven oscillations at $R_{1}=0.425g_{0}^{-1}$ and $f=120$ Hz in (a,d), to chaotic dynamics by increasing the frequency to $f=400$ Hz in (b,e), and finally to subharmonic oscillations with frequency $f/3$ by subsequently increasing the resistance to $R_{1}=1.0g_{0}^{-1}$ while maintaining $f=400$ Hz in (c,f).
To quantify the complexity of the output voltage $U_{\mathrm{out}}(t)$ in greater detail, we count the number of distinct output voltage values at discrete times $t=n/f$, where $n\in\mathbb{N}$, i.e. at time points corresponding to integer multiples of the input signal’s periodicity. The number of recurrences of distinct output values is defined as $N_{\mathrm{rec}}=\left|\{U_{\mathrm{out}}(n/f)\quad\text{for }n\in\mathbb{N}\}\right|$, where ||$\cdot$|| ⋅ | denotes the number of elements in the set. In practice, we consider two output values to be the same if their (absolute) difference is smaller than a typical tolerance, $\epsilon=0.01\,\mathrm{V}$, to account for finite numerical precision. For example, this definition yields $N_{\mathrm{rec}}=1$ for the periodic oscillation in Fig. 3(a) and $N_{\mathrm{rec}}=3$ for the subharmonic oscillation in Fig. 3(c). Similarly, $N_{\mathrm{rec}}$ is a small integer for periodic output voltages $U_{\mathrm{out}}(t)$ that are commensurate with the input signal $U_{\mathrm{in}}(t)$. However, for incommensurate or chaotic output voltages, the number of distinct output values diverges, i.e. $N_{\mathrm{rec}}\rightarrow\infty$. To ensure that the computation of $N_{\mathrm{rec}}$ remains numerically tractable, we first let the system equilibrate for 100 periods of the input signal and then analyze the next 100 periods. Thus, we restrict the analysis to $n\in[101,200]$, which bounds $N_{\mathrm{rec}}\in[1,100]$. This approach gives, for instance, $N_{\mathrm{rec}}=61$ for the case of Fig.  3(b) that we would deem to be chaotic for all intents and purposes. In the analysis and figures below, we further simplify the presentation and only distinguish $N_{\mathrm{rec}}\in[1,20]$.
Figure 4: Heat map of the number of recurrences $N_{\mathrm{rec}}$ of distinct output voltage values at times corresponding to integer periods of the input (see text), as a function of the driving frequency $f$ and the resistance of $R_{1}$ in the circuit shown in Fig. 2. Results are shown for the two positions of the switch SW1, which couples either (a) the memristors or (b) the Ohmic resistors to the RCL oscillator. All other parameter values are provided in the text. Numbers in (a) indicate $N_{\mathrm{rec}}$ for the corresponding bands. Marked points show the state-points of Fig. 3. Pink left pointing triangle for Fig. 3 (a), green upwards pointing triangle for (b) and orange right pointing triangle for (c).
In Fig. 4(a) we show the heatmap of $N_{\mathrm{rec}}$ for the circuit of Fig. 2 as a function of $R_{1}$ and $f$, with the switch SW1 in the “memristor state” and with all other parameters given in the text. We can clearly distinguish a (fully purple) low-resistor regime $g_{0}R_{1}<0.4$ where the oscillator is completely stable with $N_{\mathrm{rec}}=1$, regardless the driving frequency $f$. The high-resistor regime $g_{0}R_{1}>0.4$ shows a richer palette of behavior with large chaotic regimes (yellow) where $N_{\mathrm{rec}}\geq 20$ and substantial stable bands (purple) with $N_{\mathrm{rec}}=1$ dispersed by tiny intermediate regimes (green). Several of the stable (purple) bands extend deep into the high-resistor regime, especially at high frequencies $f>100\,\mathrm{Hz}$. Also, for $g_{0}R_{1}>10$ and low frequencies $f<10^{1.5}\,\mathrm{Hz}$, a regime appears where chaos and stability are closely inter-dispersed. Furthermore, we note that the stable (purple) bands with subharmonic response are significantly wider when the response frequency is an odd multiple of the input frequency as compared to an even multiple.
In order to analyze (and appreciate) the role of the memristors, we show the heat map of $N_{\mathrm{rec}}$ again in Fig. 4(b), however now with the switch SW1 set to the “resistor state” with two parallel Ohmic resistors replacing the two parallel memristors. The value $R_{M_{1,2}}=0.25\,\mathrm{g_{0}}^{-1}$ that we chose for the two resistors might seem a bit low at first sight, however this is justified because one of the two memristors in (a) is always in the low-resistance state due to their anti-parallel wiring. Comparing the Ohmic case of Fig. 4(b) with the memristor case of (a) reveals a similar low-resistance stability regime, however in the predominantly chaotic (yellow) regime at higher resistances much fewer stable (purple) bands appear and their (purple-yellow) interfaces appear to be sharper in (b) indicating fewer modes with intermediate harmonic periods of several driving periods. Moreover, the finely dispersed stable-and-chaotic regime found at low $f$ and high $R_{1}$ in (a) has disappeared altogether in (b). In other words, the memristors increase the regime of predictability in the predominantly chaotic high-resistance regime, and therefore contribute to a circuit being at the edge of chaos. For this reason, we focus on circuits of Memriki oscillators below.